hello,
i know i have done this sort of problem in the past but can't get answer that matches the book the answer 7 the problem is 8x - 13 = 43

Answer by pwac(253)

8x-13=43 Add 13 to both sides
8x-13+13=43+13 to give:
8x=56 Divide both sides by 8
x=7
Pete